OpenStudy (anonymous):

Suppose a+4b+9c+16d+25e=-13, 4a+9b+16c +25d +36e =-8, 9a+16b+25c+36d+49e=3, Find the value of 16a+25b+36c+49d+64e=?

OpenStudy (anonymous):

I dont think there is only one answer

OpenStudy (anonymous):

how

OpenStudy (anonymous):

we 3 equation and 5 variables

OpenStudy (anonymous):

have^

OpenStudy (anonymous):

so. we have to calculate these three equation in such a way that , we get a value of 16a+25b+36c+49d+64e=?

OpenStudy (anonymous):

But there will be many solutions

OpenStudy (anonymous):

okay..

OpenStudy (shubhamsrg):

I guess we will have to transform the 3 given eqns so that we reach the 4th required eqn.

OpenStudy (anonymous):

yup..this is vat i m trying 2 say

OpenStudy (shubhamsrg):

got it! multiply 2nd eqn by -3 and 3rd eqn by +3 and add all the 3 equations together This should help.

OpenStudy (anonymous):

x^2 -3(x+1)^2+3(x+2)^2 x^2-3x^2 -6x-3 +3x^2 + 12x+12 x^2 +6x + 9 (x+3)^2

OpenStudy (anonymous):

That works @shubhamsrg

OpenStudy (anonymous):

px^2 +q(x+1)^2 + r(x+2)^2 (p+q+r)x^2 + (2q+4r)x + (q+4r) So, p+q+r=1 ..i 2q+4r=6 ........ii q+4r=9.............iii

OpenStudy (anonymous):

Solve that three equations and multiply all three equations a+4b+9c+16d+25e=-13, 4a+9b+16c +25d +36e =-8, 9a+16b+25c+36d+49e=3, with p,q and respectively and add all of the three equations u will get the value of 16a+25b+36c+49d+64e

OpenStudy (anonymous):

(p,q and r)*

OpenStudy (anonymous):

U will get p=1 q=-3 and r=3

OpenStudy (anonymous):

thank u so much @shubhamsrg